OpenStudy (anonymous):

use the discriminant to determine the number of solutions of the quadratic equation, and whether the solutions are real or complex. x2 + 6x - 7 = 0 z2 + z + 1 = 0 (3)1/2y2 - 4y - 7(3)1/2 = 0 2x2 - 10x + 25 = 0 2x2 - 6x + 5 = 0

OpenStudy (anonymous):

for the first one: you have to identify the a,b, and c. In this case they are 1,6,-7 respectivley

myininaya (myininaya):

I will do the 4th for you ok? \[ax^2+bx+c=2x^2-10x+25=0=>a=2,b=-10,c=25\] to find the discriminant you do \[b^2-4ac=(-10)^2-4(2)(25)=100-8(25)=100-200=-100<0\] so 2 complex solutions

OpenStudy (anonymous):

now evluate: b^2-4ac

OpenStudy (anonymous):

(6)^2-4(1)(-7)=36+28=54 which means 2 real unequal roots

OpenStudy (anonymous):

i need all five because I'm clueless

OpenStudy (anonymous):

cant you use the ones we did to solve the rest?